Grade 11General Physics

a ball rolls off the edge of a horizontal table top 4m high. if it strikes the flor at apoint 5 m horizontally away from the edge of the table , what was its speed at the instant it left the table?

anirban deb
12 Years agoGrade 11

1 Answer

Apoorva Arora
12 Years ago
initial vertical velocity is zero.
So,u_{y}=0
y=u_{y}t+\frac{1}{2}gt^{2}
y=\frac{1}{2}10t^{2}
t=\sqrt{\frac{4}{5}}seconds
in the y axis, there is no acceleration.
so,a_{x}=0
x=u_{x}t+\frac{1}{2}a_{x}t^{2}
x=u_{x}t
u_{x}=\frac{5\sqrt{5}}{2}m/s
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty